Spray painter :Jaco Manetti is employed as a spray painter at NBW on a contract. Since Jaco is on contract he opted to be paid his wage on a daily basis. As such he is remunerated on a differential piecework system based on the number of panels he sprays on a day. His supervisor is responsible for the allocation of panels to be worked on a particular. The supervisor ensures that the panels allocated would ensure a fair remuneration for Jaco.

Standard time allowed: 20 minutes per panel

Standard work day: 9 hours

Normal Wage rate: N$36 per hour

Premium: 90% of piecework rate if below standard

120% of piecework rate if standard or above standard

Jaco s production for the day: 32 units

Required:

Calculate the remuneration for Jaco for the day using the differential piecework system.



1
Expert's answer
2022-03-15T12:43:35-0400

"32\times20=640" minutes


"\frac{640}{60}=10.66" hours


"10.66\times36+1.20\times10.66\times36=383.76+460.512=844.24" $